Grade 12Mechanics

An electric field of 6×10^4 volts/meter is applied on the sample of Neon gas at N.T.P calculate the dipole moment induced

fred aloyce fred
8 Years agoGrade 12

2 Answers

Arun
8 Years ago
 
Dear student
 
the dipole moment p induced in an atom of neon gas whe an electric field E is applied. , is
P = \alpha\varepsilono E
 
P = np = \varepsilono(K –1)E
 
p = \varepsilono (K –1)E / N
 
the volume of  1 m atoms of neon at N.T.P. is 22.4 litres.
 
N = Na / V
 
Hence on calculation
 
p = 2.65 * 10 ^ –36  Cm
 
 
Regards
Arun (askIITians forum expert)
Siddharth Singh
8 Years ago
the dipole moment p induced in an atom of neon gas whe an electric field E is applied. , isP = \alpha\varepsilono E P = np = \varepsilono(K –1)E p = \varepsilono (K –1)E / N the volume of 1 m atoms of neon at N.T.P. is 22.4 litres. N = Na / V Hence on calculation p = 2.65 * 10 ^ –36 Cm